### v2

# Magic-wavelength nanofiber-based two-color dipole trap with sub-$\lambda/2$ spacing

Abstract
We report on the realization and characterization of a novel magic-wavelength nanofiber-based two-color optical dipole trap for cesium that allows us to generate two diametral periodic one-dimensional arrays of trapping sites with a spacing significantly smaller than half the resonant free-space wavelength of the cesium D2-transition. This is achieved by launching a blue-detuned partial standing wave and two red-detuned light fields through the nanofiber. We trap and optically interface the atoms in the resulting periodic optical potential and characterize the trap by measuring the trap frequencies in the radial and axial direction, yielding good agreement with theoretical predictions. The implementation of this nanofiber-based optical interface with magic trapping wavelengths and sub-$\lambda/2$ spacing is an important step towards the exploration of novel collective radiative effects, such as selective radiance.
Nanofiber-based atom–light interfaces have become a powerful tool for controlling and harvesting the interaction between propagating light fields and laser trapped atoms [1, 2, 3, 4]. They allow for coupling arrays of individually trapped atoms to the nanofiber-guided mode via the evanescent field surrounding the nanofiber. This experimental platform has enabled studies and applications ranging from the efficient resonant [5, 6] and dispersive [7, 8] coupling of light and optically trapped atoms to the demonstration of propagation direction-dependent attenuation [9] and amplification [10] of light. In many of these works, the atoms could be regarded as independent emitters, e.g., contributing linearly to the total attenuation of the light field. More recently, considerable interest has been directed toward the study of collective radiative effects with nanofiber-coupled atoms mediated by the fiber-guided mode [11, 12]. This lead to the realization of optical memories [13, 14, 15] and atom-based Bragg-mirrors [16, 17] with this platform as well as the observation of sub- and superradiant emission of light, both in the single excitation sector [18, 19, 20] and with fully inverted atomic ensembles [21]. Moreover, a collective enhancement of the nonlinear response of the atoms was demonstrated allowing one to generate and tailor strongly correlated quantum states of light [22, 23, 24]. Even richer physics is expected when the interatomic distance is smaller than half the resonant emission wavelength of the atoms, $\lambda$. In this case, the atoms interact with each other via near-field dipole-dipole interactions as well as via the common nanofiber-guided mode. An emblematic effect predicted in this regime is selective radiance, where the scattering of guided light back into the waveguide mode is collectively enhanced while it is suppressed for all unguided modes. This would enable low-loss propagation of light past the atomic array and may allow one to realize quantum devices such as quantum memories with significantly enhanced performance [25, 26, 27, 28, 29, 30]. Here, we demonstrate a nanofiber-based two-color dipole trap with distance $d\simeq 0.35\lambda$ between adjacent trapping sites, i.e., significantly smaller than $\lambda/2$. We use magic wavelengths for the trapping fields in order to minimize the inhomogeneous broadening of the optical transition frequency due to trap-induced Stark shifts, thereby providing optimal conditions for the implementation and observation of novel collective radiative effects.
The trapping scheme is based on a two-color nanofiber-guided dipole trap [31] realized by launching laser fields, which are red- and blue-detuned to the D2-transition of cesium, through a tapered optical fiber with a nanofiber-waist of $a=200\,\mathrm{nm}$ radius. The blue and red-detuned laser fields are launched from both ends into the nanofiber (see Fig. 1) and are tuned to the “magic” wavelengths of $685\,\mathrm{nm}$ and $935\,\mathrm{nm}$, respectively, so as to suppress the differential scalar light shifts between the ground and excited states [6]. All fields are quasi-linearly polarized along the $x$-direction. The interference of the two counter-propagating blue-detuned trapping fields gives rise to a partial standing wave in the evanescent field, resulting in a dipole potential in the axial $z$-direction. Additionally, we launch two counter-propagating red-detuned laser fields through the nanofiber to create an attractive dipole potential, thereby confining the atoms in all directions, radially, axially, and azimuthally. The two counter-propagating red-detuned laser fields are detuned from each other by ∼$\sim 150\,\mathrm{GHz}$ to prevent them from forming a standing wave pattern while at the same time minimizing the inhomogeneous Zeeman broadening due to the vector light shift [6]. We note that the vector polarizability of cesium for the blue-detuned magic wavelength is two orders of magnitude smaller than that for the red-detuned magic wavelength [32]. Therefore, the vector light shift for the blue-detuned dipole trap is negligible. Figure 2 a) shows the total trapping potential, $U_{\mathrm{tot}}$, which is the sum of the blue and red-detuned dipole potentials as well as the van der Waals potential induced by the fiber surface [31]. The intensities of the two red-detuned fields are equal at the position of the atoms. Each of the two counter-propagating beams has a total transmitted power of $1\,\mathrm{mW}$. In order to provide a potential barrier that prevents atoms from escaping towards the fiber surface in the nodes of the blue-detuned standing wave, we use unbalanced powers of the two counter-propagating blue-detuned guided fields of $P_{\mathrm{blue,1}}=0.6\,\mathrm{mW}$ and of $P_{\mathrm{blue,2}}=16\,\mathrm{mW}$, respectively. As shown in Fig. 2 b) and c), this configuration creates two diametral one-dimensional arrays of trapping sites along the $z$-direction, where adjacent trapping sites are separated by $d=\lambda_{\mathrm{blue}}/(2n)=300\,\mathrm{nm}$. Here, $n=1.14$ is the optical mode’s effective refractive index. In the radial direction, the atoms are located around the local potential minimum at a distance of about 350 nm from the fiber surface, see Fig. 2 a). Finally, in the azimuthal $\varphi$-direction, the quasi-linear polarization of both red- and blue-detuned trapping fields leads to a modulation of the evanescent field intensity around the nanofiber, which results in the confinement of the atoms in the azimuthal direction [33, 5]. For our trap configuration and parameters, the potential depth of our trap is approximately $110\mu K$. Due to the small trapping volume, our traps operate in the collisional blockade regime such that each trapping site contains at most one atom [34, 35].
Figure 1: Experimental setup of the nanofiber-based two-color dipole trap: The interference of two blue-detuned ($\lambda_{\mathrm{blue}}=685\,\mathrm{nm}$) counter-propagating light fields creates a repulsive partial standing wave dipole potential along the $z$-direction. Together with two attractive red-detuned running wave dipole potentials ($\lambda_{\mathrm{red}}=935\,\mathrm{nm}$), which are $150\,\mathrm{GHz}$ detuned from each other in order to minimize the vector light shift, we create two diametral one-dimensional arrays of trapping sites in the evanescent field of the nanofiber. One acousto-optic modulator (AOM) in the path of the blue-detuned light is used for intensity modulation and an electro-optic modulator (EOM) is used for phase modulation of the high-power arm of the counter-propagating blue-detuned trapping fields. By means of these modulations, we can study the modulation frequency dependent heating of the atoms via loss spectroscopy by probing the atomic ensemble with a laser field resonant with the D2-line at a wavelength of $852\,\mathrm{nm}$.
Figure 2: Calculated trapping potential: a) Radial dipole potential at a minimum of the blue-detuned partial standing wave potential, see dashed-white line in b). The total potential (black line) is the sum of the attractive potential created by the red-detuned trapping fields (red line), the repulsive potential created by the blue-detuned light fields (blue line) and the attractive van der Waals potential (green dashed-dotted line). b) Trapping potential in the $x$–$z$-plane, i.e., along the nanofiber, as a function of the radial distance from the nanofiber center. c) Trap potential in the $x$–$y$-plane, where the $z$-position corresponds to a minimum of the blue-detuned standing wave potential. The potential is calculated for powers of $0.6\,\mathrm{mW}$ and $16\,\mathrm{mW}$ for the blue-detuned trapping fields and equal powers of $2\times 0.5$ mW for the red-detuned trapping fields. All trapping fields are quasi-linearly polarized along the $x$-axis, also indicated by the yellow arrow.
In order to load atoms into the two arrays of trapping sites, we first prepare a cloud of cold cesium atoms by using a magneto-optical trap. This cloud overlaps with a section of the nanofiber, and in a subsequent molasses cooling stage the atoms are loaded into the nanofiber-based dipole trap. In order to estimate the number of loaded atoms, we measure the optical depth, (OD), of the trapped ensemble. To do this, we switch off the cooling laser and interrogate the atomic ensemble by launching a probe laser field through the nanofiber. We scan the laser frequency across the $56S_{1/2}\;F=4\rightarrow 6P_{3/2}\;F^{\prime}=$ 5 D2-transition of the cesium. Afterward, we determine the OD of the ensemble by fitting the transmission spectrum with a saturated Lorentzian function. We obtain an optical depth of $\mathrm{OD}=42\pm 2$, about $3\,\mathrm{ms}$ after switching off the optical molasses cooling. We note that the probe power is small enough to neglect saturation effects. The optical depth relates to the number of trapped atoms, $N$, via $\mathrm{OD}=4\beta N$. Here, $\beta=\Gamma_{\textsc{wg}}/\Gamma$ is the atom-light coupling constant, which describes the ratio of the spontaneous emission rate into the forward direction of the waveguide, $\Gamma_{\textsc{wg}}$, and the total emission rate $\Gamma$. Assuming a typical coupling factor $\beta\approx 1\%$ for our radial trapping distance [23], the total number of trapped atoms is on the order of $N\approx 1000$. In order to determine the lifetime of the atoms in the trap, we measure the OD as a function of the delay, $t$, between loading and the center of the $1\,\mathrm{ms}$ probe pulse. We obtain the data shown in Fig. 3 a). Fitting an exponential decay of the OD, we find a trap lifetime of $\tau=8.5\,\mathrm{ms}\pm 0.4\,\mathrm{ms}$. In order to confirm that the measured transmission signal is indeed due to trapped atoms rather than untrapped atoms from the residual atomic cloud surrounding the nanofiber, we carry out an additional measurement. Its experimental sequence is similar to the one used for the above lifetime measurement, with the difference that at $t=3.4\,\mathrm{ms}$, see gray dashed line in Fig. 3 b), we switch off the red-detuned trapping field for a duration $\mathrm{\Delta T_{off}}$. After a duration $\mathrm{\Delta T_{off}}$ of 3.5 $\mu$s (cyan), 350 $\mu$s (green) or 3500 $\mu$s (orange), we switch the red-detuned trapping field on and measure the remaining OD. As shown in Fig. 3 b), we observe a sharp drop of the OD when the red-detuned laser is switched off. This indicates that the atoms were indeed trapped before the switch-off because the latter does not affect the density of the residual atomic cloud surrounding the nanofiber. For the shortest switch-off time, we observe a higher residual OD than in the other two cases. This can be explained by the fact that $\mathrm{\Delta T_{off}}=$ 3.5 $\mu$s is comparable with the trap oscillation period, see below, such that some atoms are recaptured in the dipole trap when switching the trapping field back on.
Figure 3: Measurement of the OD of the ensemble of trapped atoms for different hold times after switching of the cooling laser: a) The exponential fit to the data yields a lifetime of $\tau=8.5\,\mathrm{ms}\pm 0.4\,\mathrm{ms}$. b) The red-detuned trapping light field is switched off at $t=3.4\,\mathrm{ms}$ (dotted gray line). After $\Delta T_{\mathrm{off}}$ of 3.5 $\mu$s (cyan), 350 $\mu$s (green), or 3500 $\mu$s (orange) the red-detuned trapping field is switched on again. A significant drop of the OD (colored) compared to the regular lifetime (black) is visible, confirming that most of the interrogated atoms during the lifetime measurement are indeed trapped. Experimental parameters are as in Fig. 2. The error bars indicate the $1\sigma$ standard deviation of the measured OD.
Figure 4: Measurement of the trap frequencies: Measured OD when modulating a) the intensity of both blue-detuned trapping fields and b) the phase of one of the two. The OD is determined by a fit of the transmission spectra after heating the atoms in the trap configuration as shown in Fig. 2 ($P_{\mathrm{red,tot}}=1\,\mathrm{mW}$). We identify the resonance frequencies by performing a moving average on the raw data (gray) resulting in the black curves, and then by searching for local minima. The determined local minima for the intensity modulation (black squares) and phase modulation (black circles) can then be compared to the calculated fundamental trap frequencies (vertical solid lines) and the second harmonics of the trap frequencies (vertical dashed lines) for the radial ($f_{r}$) and axial ($f_{z}$) motion. The step size of the applied modulation frequency for the intensity modulation is $10.2\,\mathrm{kHz}$, which we here take as an estimate of the uncertainty of the measured trap frequencies, yielding an error bar of ±$\pm 5.1\,\mathrm{kHz}$. Likewise, the error bar for the phase modulation is estimated to be ±$\pm 2.5\,\mathrm{kHz}$. c) Trap frequencies obtained from measurements identical to those shown in panels a) and b) for different powers of the red-detuned laser. We observe good agreement of corresponding theory predictions (solid and dashed lines) with the determined resonance frequencies for the intensity modulation (black squares) and for the phase modulation (black circles).
In order to characterize the trap, we measure the trap frequencies along the radial ($f_{r}$) and axial ($f_{z}$) direction. To do this, we perform loss spectroscopy based on a resonant heating technique [36] using an acoustic-optic modulator (AOM) and an electro-optical modulator (EOM) as depicted in Fig. 1. First, we modulate the intensity of the blue-detuned trapping field with the AOM before it is split and launched in the two counter-propagating directions through the nanofiber. The modulation of the intensity translates into a modulation of the center position of the trapping potential in the radial direction. After loading the trap, we modulate the potential for 2 ms and then measure the OD. When the modulation frequency $f_{\text{mod}}$ reaches a harmonic of the trap frequencies, atoms are heated out of the trap and the OD decreases. To identify the resonance frequencies for which the atoms are subject to maximal heating, we first perform a moving average on the OD extracted from the transmission data to mitigate the effect of shot-to-shot fluctuations of the number of trapped atoms. Afterwards, we apply a search for local minima algorithm on the smoothed data set. The results of the measurement are shown in Fig. 4 a). We observe local minima of the OD at $116\,\mathrm{kHz}$ and $232\,\mathrm{kHz}$ (black squares). The measured resonances agree well with the calculated fundamental and second harmonic radial trap frequencies at $109\,\mathrm{kHz}$ and $218\,\mathrm{kHz}$ (red vertical lines). In order to measure the axial trap frequency ($f_{z}$), we use an EOM that modulates the phase of one of the two blue-detuned trapping fields, see Fig. 1. The EOM is aligned such that it only modulates the phase of the corresponding blue-detuned trapping field without affecting its polarization. This phase modulation, applied for a duration of 5 ms, causes a sinusoidal modulation of the position of the standing wave potential along the $z$-direction. Figure 4 b) shows the measured OD as a function of the modulation frequency. We observe two local minima at $139\,\mathrm{kHz}$ and at $235\,\mathrm{kHz}$ (black circles) that are close to the calculated fundamental axial trap frequency of $134\,\mathrm{kHz}$ (blue vertical line) and the second harmonic of the radial frequency at $218\,\mathrm{kHz}$ (red dashed vertical line). In order to perform a systematic characterization of the new trap configuration, we repeat the measurement of the trap frequencies for different powers of the red-detuned trapping field. The resulting trap frequencies are shown in Fig. 4 c). We observe good agreement of the measured trap frequencies with the calculated fundamental frequencies (solid lines) and second harmonic (dashed line). As apparent in Fig. 4 b) and c), the phase modulation also excites the second harmonic radial trap frequency and not only the fundamental axial trap frequency. We attribute this effect to the anharmonicity of the trapping potential, see Fig. 2, which leads to a coupling between, e.g., the axial and radial degree of freedom of the atoms [37]. This explains why we also observe parametric heating in the radial direction at $2f_{r}$ [38].
In summary, we have experimentally demonstrated magic-wavelength trapping of laser-cooled atoms in a novel type of nanofiber-based two-color dipole trap. By using a blue-detuned partial standing wave and two red-detuned running waves, we create two diametral one-dimensional arrays of trapping sites for cesium atoms. In this trapping potential, the spacing of two adjacent sites is $d\simeq 0.35\lambda$, i.e. only 35 % of the resonant free-space wavelength of the cesium D2 line. We determine the lifetime of the atoms in this new trapping configuration to be $\tau=8.5\,\mathrm{ms}\pm 0.4\,\mathrm{ms}$ and perform a systematic characterization of the radial and axial trap frequencies which agree well with the expectations. This realization and characterization of a sub-$\lambda/2$ spaced standing wave nanofiber-guided dipole trap is an important step towards the experimental study of novel collective radiative effects, such as selective radiance in 1D atomic arrays.
Acknowledgment
M. C. and M. S. acknowledge support from the European Commission (Marie Skłodowska-Curie IF Grants Nos. 101029304 and IF Grant No. 896957). We acknowledge funding by the Alexander von Humboldt Foundation in the framework of the Alexander von Humboldt Professorship endowed by the Federal Ministry of Education and Research, funding by the European Commission under the project DAALI (No. 899275), and by the Einstein Foundation (Einstein Research Unit on Quantum Devices). We thank Christoph Clausen, Christian Liedl, and Sebastian Pucher for developing and improving the trap potential calculations.